OpenStudy (hockeychick23):

Of 600 students sampled, 480 said they hoped to own a house someday. With 95% confidence, what is the approximate percentage of the students in the population who hope to own a house someday?

OpenStudy (haseeb96):

480/600 x. 100 Now solvsolve it u will get your answer

OpenStudy (hockeychick23):

80%

OpenStudy (hockeychick23):

its supposed to be a range tho either: A. 73.5% to 86.5% B. 78.4% to 81.6% C. 75.1% to 84.9% D. 76.7% to 83.3%

OpenStudy (anonymous):

I think it might be A. I think the distribution is binomial because either a student wants to own a house or doesn't. The mean of the distribution is 480, or p = 80%. The variance of a binomial distribution is np(1-p). np(1-p) = 600*0.8*(1-0.8)=96. Variance is standard deviation squared, so standard deviation is square root of 96, or about 19.6. There's this approximation: http://en.wikipedia.org/wiki/68%E2%80%9395%E2%80%9399.7_rule Which says that 95% of values lie within +/-2 standard deviations. So 480 + 19.6*2 and 480-19.6*2, which 519.2 and 440.8. Divide by 600 to get percentages 73.5% to 86.5%.

OpenStudy (anonymous):

I could be completely wrong. But some reasoning is better than a random guess.